In carbon monoxide molecule d carbon atom n d oxyen r seprated by dist 1.2 angstrom the dist of c.o.m frm d carbom atom is

Grade:11

2 Answers

Tarun Sethi
6 years ago
The Carbon Oxygen bond behave like a weight with Carbon and Oxygen at two ends. Now the distance of carbon atom is directly proportional to the mass of carbon atom. Let us consider distance from carbon atom to be `x` Angstrom so distance of oxygen atom will be (1.2 - x) Angstrom. Therfore, 12 (mass of carbon atom) ÷ 16 (mass of oxygen atom)= (x) ÷ (1.2-x)After calculation we get x = 0.685 Angstrom
Tarun Sethi
6 years ago
Dry I did a little mistake above the mass is inversely proportional to distance of atom from centre of mass so12 (mass of carbon atom) ÷ 16 (mass of oxygen atom)= (1.2-x) ÷ (x)After calculation we get x = 0.685 Angstrom